The smart treasure fund was credited for Samson after he lost his leg in a battle with Pirates the fund was undertaken to pay him 1200000 now Samson prefers to receive three payments one 3 years from now: one twice the size of the first payment 6 years from now, and one four times the size of the first payment 10 years from now. The amount of money to the nearest rent that Samson can expect to receive 6 years from now if the interest rate applicable is 8.6% per year compounded quarterly will be?


1
Expert's answer
2022-03-07T02:35:02-0500

A=P(1+r/n)nt

A-final amount

P-initial principal

r-rate of interest

n-number of times interest is compounded per time period

t- number of periods

In this case the first payment is the determinant of all the other payments

1st payment

A=?

P-1200000

r-8.6%

n-4

t-3

A=1200000(1+0.86/4)4×3

=1200000(1.215)12

=1200000×10.35

First payment =12,419,284-1,200,000

=11,219,284

6 years equals twice to first payment =22,438,568.30